OpenStudy (anonymous):

PLease im being timed this needs to be in simplified form. x-7/2x-12+6/x^2-36

OpenStudy (anonymous):

Is it 7 x - ---- x - 12 + 6 2 -------------------- ??? x^2 - 36

OpenStudy (anonymous):

x-7 over 2x-12 plus 12+6/x^2-36

OpenStudy (anonymous):

\[\frac{x+5}{2x+12}\]

OpenStudy (anonymous):

is your answer. it wants you to add

OpenStudy (anonymous):

x - 7 12 + 6 -------- + --------- ??? 2x - 12 x^2 - 36

OpenStudy (anonymous):

yes harkirat!

OpenStudy (anonymous):

i am assuming you meant \[\frac{x-7}{2x-12}+\frac{6}{x^2-36}\]

OpenStudy (anonymous):

because no one writes \[12+6\] when they could just write 18

OpenStudy (anonymous):

oh ur right its just 6

OpenStudy (anonymous):

that is what i thought. my answer is right

OpenStudy (anonymous):

yes u are right sorry

OpenStudy (anonymous):

x - 7 12 + 6 -------- + --------- 2x - 12 x^2 - 36 x - 7 12 + 6 -------- + ----------- 2(x - 6) (x+6)(x-6) (x-7)(x+6) + 2(18) --------------------- 2(x-6)(x+6) x^2 - x - 42 + 36 x^2 - x - 6 x^2 -3x + 2x - 6 ---------------- = -------------- = ---------------- 2(x-6)(x+6) 2(x-6)(x+6) 2(x-6)(x+6) x(x-3) + 2(x-3) (x+2)(x-3) -------------- = ------------ 2(x-6)(x+6) 2(x-6)(x+6)